OpenStudy (anonymous):

A wire of length L meters is to be divided into two parts; one part will be bent into a square and the other into a circle. How should the wire be divided to make the sum of the areas of the square and circle as small as possible?

OpenStudy (anonymous):

i hate this problem but you are persistent so we can do it i think

OpenStudy (anonymous):

lol hey I figured out the maximizing part...

OpenStudy (anonymous):

put \(x\) as the amount for the circle, then \(L-x\) is the amount for the square. circle will have area ... you did?? then your are done

OpenStudy (anonymous):

no this is minimizing, the maximizing part was a trick question. It was zero for the square and L for the circle

OpenStudy (anonymous):

Infinitely small circle, rest is square

OpenStudy (anonymous):

oh i doubt that

OpenStudy (anonymous):

but how do I write that

OpenStudy (anonymous):

oooh minimize the area???

OpenStudy (anonymous):

Since the area of a circle is always larger than a square for the same perimeter, you would want to minimize the circle

OpenStudy (anonymous):

sorry i read it incorrectly @seattle_ice i think has it, let me shut up

OpenStudy (anonymous):

what did you get for the total area in terms of \(x\) ?

OpenStudy (anonymous):

(9x-L)/8

OpenStudy (anonymous):

wait that was the derivative, hold on

OpenStudy (anonymous):

(x^2)/16 + pi* [ (L-x)/2pi ]^2

OpenStudy (anonymous):

hmm i got something different area of the square is \((\frac{L-x}{4})^2\) and area of the circle is \(\frac{x^2}{4\pi}\) i think

OpenStudy (anonymous):

oh i see i switched from what you had

OpenStudy (anonymous):

are you talking about the total area or just the circle?

OpenStudy (anonymous):

no matter it should work out the same either way

OpenStudy (anonymous):

i used \(L-x\) for the perimeter of the square, and \(x\) for the circumference of the circle

OpenStudy (anonymous):

Oh, yeah I did the opposite

OpenStudy (anonymous):

thought it might be easier having a simpler expression with the \(\pi\) and all

OpenStudy (anonymous):

no matter it should work out the same

OpenStudy (anonymous):

alright, but I'm still not sure where I went wrong

OpenStudy (anonymous):

you get \[A(x)=\frac{x^2}{16}+\frac{(L-x)^2}{4\pi}\] in your method

OpenStudy (anonymous):

derivative would be \[A'(x)=\frac{4x+\pi x-4L}{8\pi}\]

OpenStudy (anonymous):

if you want the critical point, i guess it would be better to write \[A'(x)=\frac{(4+\pi) x-4L}{8\pi}\] then set \[(4+\pi)x-4L=0\] and solve for \(x\) in two easy steps

OpenStudy (anonymous):

wow, I guess my algebra was wrong

OpenStudy (anonymous):

wow, thanks so much! :P really, it would have taken me w while to find my error!

OpenStudy (anonymous):

it would be a lot easier to use \(x\) as the circumference of the circle and \(L-x\) as the perimeter of the square

OpenStudy (anonymous):

that way you get \[A(x)=(\frac{L-x}{4})^2+\frac{x^2}{4\pi}\] and taking the derivative is easier

OpenStudy (anonymous):

as for the max, you have only one critical point, max occurs there, and min will occur at one of the endpoints of the interval, namely \(x=0\) or \(x=L\)

OpenStudy (anonymous):

right, I'll try to rmemeber that for next time. I realize that setting up the problem can be more importatn than actually doing the problem..

OpenStudy (anonymous):

yeah good luck pretty sure min is make all a square, max is some point in the middle